OpenStudy (anonymous):

Help? Please simplify: 2-x/3 ----- 5+x/8

OpenStudy (anonymous):

\[=\frac{6-x}{3} / \frac{40+x}{8}\] \[=\frac{(6-x)}{3} * \frac{8}{(40+x)}\] \[=\frac{8(6-x)}{3(40+x)}\]

OpenStudy (anonymous):

Whoa! You typing a book :) I really appreciate your help.